What is the variance of the discrete random variable?


1
Expert's answer
2021-02-24T13:51:14-0500

"\text{variance of the discrete random variable the difference between}\newline
\text{ the values of a random variable and its mathematical expectation}\newline
Var(X)=\sum_{i=1}^{n}{x_i^2 \cdot p_i}-\left(\sum_{i=1}^{n}{x_i \cdot p_i} \right)^2"